OpenStudy (anonymous):

How do I find the equation of a quadratic function with points at (-4,4) & vertex (-2,-4)? f(x)=?

OpenStudy (anonymous):

We know the form y=a(x-h)^2+k (h,k) or (-2,-4) is the vertex, and since we have another point (x,y) or (-4,4) we can plug those in and solve for a. 4=a(-4+2)^2-4

OpenStudy (anonymous):

What do you get when you solve for a?

OpenStudy (anonymous):

4=4a-4 8=4a 2=a

OpenStudy (anonymous):

a=2

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

Yes, now put that in. y=2(x+2)^2-4

OpenStudy (anonymous):

If it wants standard form then multiply all of it out

OpenStudy (anonymous):

ok another example. \[y=a(x-h)^2+k ... \] Vertex (3,3) & point (5,11) \[11=a(5-3)^2+3\] \[11=a(2)^2+3\] \[(8=4a)/4\] \[a=2\] So \[y=2(x-3)^2+3\] Expanded \[y=2x^2-36+3 ... becomes ... y=2x^2-33\]

OpenStudy (anonymous):

ope ... \[y=2x^2-18+3\] ... becomes ... \[y=2x^2-15\]

OpenStudy (anonymous):

i think you lost an x in there

OpenStudy (anonymous):

when expanding

OpenStudy (anonymous):

you should get 2x^2-12x+21

OpenStudy (anonymous):

... been a long time ... where did I go wrong?

OpenStudy (anonymous):

y=2(x−3)^2+3 first do (x-3)^2 y=2(x^2-6x+9)+3 distribute the 2 y=2x^2-12x+18+3 add the 3 y=2x^2-12x+21